If f is a linear function,
f(0.1) = 9.3,
and
f(0.4) = −6.3,
find an equation
f(x)
for the function.
f(x) = _______

Answer by ewatrrr(24785)

f(0.1) = 9.3 and f(0.4) = -6.3
y = ax + b
a(.1) + b = 9.3
a(.4) + b = -6.3
|
-a(.1) - b = -9.3
a(.4) + b = -6.3
.3a = -15.6
a = 52
y = 52x + b
Using: f(0.1) = 9.3
9.3 = 5.2 + b
4.1 = b
|
y = 52x + 4.1
0r
f(x) = 52x + 4.1